# A note on the scatteredness of reflection orders

### v1

@@ 2.5. Normalized roots
Assume that $V=\mathbb{R}\Pi$. For $x\in V$, we denote by $(x_{1},x_{2},\dots,x_{n})$ the coordinates of $x$ with respect to the basis $\alpha_{1},\alpha_{2},\dots,\alpha_{n},\alpha_{i}\in\Pi$. Consider the affine hyperplane $P$ in $\mathbb{R}^{n}:x_{1}+x_{2}+\cdots+x_{n}=1$. We call $P$ the standard affine hyperplane in $\mathbb{R}^{n}$. The set of the normalized roots of $(W,S)$, denoted by $\widehat{\Phi}$, is the intersection of $\bigcup_{\beta\in\Phi}\mathbb{R}\beta$ with $P$. For any $\beta=\sum_{i=1}^{n}k_{i}\alpha_{i}\in\Phi^{+}$, there is a unique normalized root $\widehat{\beta}=\frac{\beta}{\sum_{i=1}^{n}k_{i}}\in\widehat{\Phi}$ associated with it. We call
$(\frac{k_{1}}{\sum_{i=1}^{n}k_{i}},\frac{k_{2}}{\sum_{i=1}^{n}k_{i}},\cdots,\frac{k_{n}}{\sum_{i=1}^{n}k_{i}})$
the barycentric coordinates of $\beta$.
@@ 2.6. Lexicographic reflection order
Let $V=\mathbb{R}\Phi$ and let the root basis $\Pi=\{\alpha_{1},\alpha_{2},\dots,\alpha_{n}\}$. Choose an ordered basis $\{v_{1},v_{2},\cdots,v_{n}\}$ of $V$. A natural choice of this basis is often $v_{i}=\alpha_{i},1\leq i\leq n$ if the root basis is linearly independent, but we do not assume this. Each root can be written uniquely as $\sum_{i=1}^{n}k_{i}v_{i}.$ This induces a lexicographic ordering on $\mathbb{R}^{n}$:
@@ 2.8. Limit root and isotropic cone
The normalized isotropic cone is the intersection of the standard affine hyperplane with the isotropic cone. An accumulation point of $\widehat{\Phi}$ (the set of projective roots) is called a limit root. It is proved in [10] that all limit roots lie on the normalized isotropic cone. The concept of limit roots was introduced recently as a tool for better visualizing the positive roots. It has proved instrumental in investigating the distribution of roots in infinite Coxeter systems of low rank, yielding deeper insight into concepts such as the imaginary cone, biclosed sets, the dominance order, and reflection subgroups. For this notion and its application, see [4], [6], [8], [10], [9], [11] and [13].
@@
Figure 1. The normalized roots and the normalized isotropic cone of rank 3 universal Coxeter system
@@ 3. Dense Subset
Let $(W,S)$ be a rank 3 universal Coxeter system, i.e. $S=\{s_{1},s_{2},s_{3}\}$ and
@@
Let the root system $\Phi$ be constructed via the standard reflection representation as in Subsection 2.2. Impose a total order on $\Pi$ by requiring $\alpha_{s_{1}}\prec\alpha_{s_{2}}\prec\alpha_{s_{3}}$. Denote the associated lexicographic reflection order by $\prec_{\mathrm{reflex}}$.
@@
is a positive root. Therefore $c_{1}+2c_{2}+2c_{3}\geq 0$ and consequently $c_{2}+c_{3}\geq\frac{c_{1}}{2}$. If $c_{1}>0$, then
@@
Assume that $c\in(0,\frac{2}{3})$ and that the set
@@
is infinite. Furthermore this set is precisely the set of the positive roots of a maximal infinite dihedral reflection subgroup.
@@
contains at least two elements. There exists a unique maximal dihedral reflection subgroup $W^{\prime}$ whose set of positive roots $\Phi_{W^{\prime}}^{+}$ contains these two roots. Furthermore $\Phi_{W^{\prime}}^{+}$ consists precisely of those positive roots whose first barycentric coordinate is $c$, and the normalized roots of them lie on a line segment parallel to the side connecting $\widehat{\alpha_{2}}$ and $\widehat{\alpha_{3}}$ of the equilateral triangle in Figure 1 (as depicted by the green line segment there). Since $c<\frac{2}{3}$, this line segment intersects with the normalized isotropic cone (the circle inscribed in the triangle). By Proposition 1.5(ii) in [10] this implies that $W^{\prime}$ is an infinite dihedral reflection subgroup and therefore $\Phi_{W^{\prime}}^{+}$ is also infinite. Hence, there are an infinite number of positive roots with $\frac{c_{1}}{c_{1}+c_{2}+c_{3}}=c$.
If otherwise $c_{2}=c_{3}=k$, we have $B(c_{1}\alpha_{1}+k\alpha_{2}+k\alpha_{3},c_{1}\alpha_{1}+k\alpha_{2}+k\alpha_{3})=1.$ Therefore, $c_{1}^{2}-4c_{1}k=c_{1}(c_{1}-4k)=1$. Since $c_{1}$ and $c_{1}-4k$ are both integers, $c_{1}=1,k=0$ and thus $c_{1}\not\in(0,\frac{2}{3})$. Therefore the assertion of the Lemma follows. ∎
@@
The set $\{\frac{c_{1}}{c_{1}+c_{2}+c_{3}}\mid\,c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\}\cap(0,\frac{2}{3})$ is dense in the interval $(0,\frac{2}{3})$.
@@
Let $x,y\in[0,\frac{2}{3}],x\neq y$ be such that there exist two positive roots $\alpha,\beta$, with first barycentric coordinates $x$ and $y$ respectively.(It is possible to find such $x,y$ as one can take 0 and $\frac{2}{3}$.) By Lemma 3.2, there exists a maximal infinite dihedral reflection subgroup $W^{\prime}$ (resp. $W^{\prime\prime}$) whose set of positive roots contains $\alpha$ (resp. $\beta$). Furthermore, any positive root of $W^{\prime}$ (resp. $W^{\prime\prime}$) has the first barycentric coordinate equal to $x$ (resp. $y$). Since the limit roots of $W^{\prime}$ and $W^{\prime\prime}$ lie on the (normalized) isotropic cone, two roots $\alpha^{\prime}\in\Phi_{W^{\prime}}^{+},\beta^{\prime}\in\Phi_{W^{\prime\prime}}^{+}$ can be chosen such that $\widehat{\alpha^{\prime}}$ and $\widehat{\beta^{\prime}}$ are close enough to the normalized isotropic cone for the line segment connecting them to cut the normalized isotropic cone. (See Figure 2 for an illustration of one possible situation.) Then by Proposition 1.5(ii) in [10], the maximal dihedral reflection subgroup containing the reflections in $\alpha^{\prime}$ and $\beta^{\prime}$ is infinite and there exists some $\gamma$ in the set of positive roots of this maximal dihedral reflection subgroup such that $\gamma$ is a positive linear combination of $\alpha^{\prime}$ and $\beta^{\prime}$. Then the first barycentric coordinate of $\gamma$ is necessarily between $x$ and $y$. ∎
@@
It is a natural question whether $\{\frac{c_{1}}{c_{1}+c_{2}+c_{3}}\mid\,c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\}\cap(0,\frac{2}{3})=\mathbb{Q}\cap(0,\frac{2}{3}).$ If not, is there a nice characterization of the set $\{\frac{c_{1}}{c_{1}+c_{2}+c_{3}}\mid\,c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\}\cap(0,\frac{2}{3})$?
@@
If $\frac{c_{1}}{c_{1}+c_{2}+c_{3}}=0,$ $\alpha:=c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}$ is a root contained in the root system of the (infinite dihedral) parabolic subgroup generated by $s_{2},s_{3}$. These roots form an initial interval of $\prec_{\mathrm{reflex}}.$ The corresponding normalized roots lie on the side connecting $\widehat{\alpha_{2}}$ and $\widehat{\alpha_{3}}$ in the equilateral triangle in Figure 1. Write $\Phi^{+}_{1}=\{c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\mid\,c_{1}=0\}.$
If $\frac{c_{1}}{c_{1}+c_{2}+c_{3}}=\frac{2}{3}$, $\alpha:=c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}$ is a root contained in the root system of the infinite maximal dihedral reflection subgroup generated by $s_{2\alpha_{1}+\alpha_{2}},s_{2\alpha_{1}+\alpha_{3}}$. The corresponding normalized roots lie on the blue line segment in Figure 1. Write $\Phi^{+}_{3}=\{c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\mid\,\frac{c_{1}}{c_{1}+c_{2}+c_{3}}=\frac{2}{3}\}.$
Write $\Phi^{+}_{2}=\{c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\mid\,\frac{c_{1}}{c_{1}+c_{2}+c_{3}}\in(0,\frac{2}{3})\}.$
@@
Note that by the previous analysis the restriction of $\prec_{\mathrm{reflex}}$ to $\Phi^{+}_{1}$ (resp. $\Phi^{+}_{3}$) has order type $\omega+\omega^{*}$.
Let $c^{\prime}\in\{\frac{c_{1}}{c_{1}+c_{2}+c_{3}}\mid\,c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\}\cap(0,\frac{2}{3})$. By Lemma 3.2, the roots in the set $C^{\prime}:=\{c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\mid\frac{c_{1}}{c_{1}+c_{2}+c_{3}}=c^{\prime}\}$ are the positive roots of an infinite dihedral subgroup. We can therefore conclude that the roots in $C^{\prime}$ form an interval $I_{c^{\prime}}$ in $\prec_{\mathrm{reflex}}$ of order type $\omega+\omega^{*}$ by the definition of lexicographic reflection order and the fact that $\omega+\omega^{*}$ is the only possible order type of a reflection order on a dihedral group. Again by the definition of lexicographic reflection order, one has $I_{c^{\prime}}\prec I_{c^{\prime\prime}}$ if and only if $c^{\prime}<c^{\prime\prime}$. On the other hand, $\Phi^{+}_{3}$ is the disjoint union of $I_{c^{\prime}},c^{\prime}\in\{\frac{c_{1}}{c_{1}+c_{2}+c_{3}}\mid\,c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\}\cap(0,\frac{2}{3})$, which is dense in $(0,\frac{2}{3})$ by Lemma 3.3. Therefore, by Theorem 2.8 of [14] we conclude that the natural total order on $\{\frac{c_{1}}{c_{1}+c_{2}+c_{3}}\mid\,c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\}\cap(0,\frac{2}{3})$ has the order type of the set of rational numbers, i.e. $\eta$ and that the restriction of $\prec_{\mathrm{reflex}}$ on $\Phi_{3}^{+}$ has the order type $(\omega+\omega^{*})\eta$.
@@
We record the following lemma for later use.
@@
is not empty. By symmetry, Lemma 3.2 guarantees that the set
@@
is precisely the set of the positive roots of a maximal infinite dihedral reflection subgroup $U^{\prime}$. The normalized roots of them lie on a line segment parallel to the side connecting $\widehat{\alpha_{1}}$ and $\widehat{\alpha_{3}}$ of the equilateral triangle in Figure 1 (as depicted by the purple line segment there). The first barycentric coordinates of the positive roots of $\Phi_{U^{\prime}}^{+}$ vary between 0 and $\frac{2}{3}$. Take $\alpha,\beta\in\Phi_{U^{\prime}}^{+}$ with the first barycentric coordinate of $\alpha$ smaller than that of $\beta$. Then $\alpha\prec_{\mathrm{reflex}}\beta.$ The by Lemma 3.3, the interval $(\alpha,\beta)$ is infinite as there are infinite many roots with first barycentric coordinate between the barycentric coordinate of $\alpha$ and the barycentric coordinate of $\beta.$

### v2

@@ 2.5. Normalized roots
Assume that $V=\mathbb{R}\Pi$. For $x\in V$, we denote by $(x_{1},x_{2},\dots,x_{n})$ the coordinates of $x$ with respect to the basis $\alpha_{1},\alpha_{2},\dots,\alpha_{n},\alpha_{i}\in\Pi$. Consider the affine hyperplane $P$ in $\mathbb{R}^{n}:x_{1}+x_{2}+\cdots+x_{n}=1$. We call $P$ the standard affine hyperplane in $\mathbb{R}^{n}$. The set of the normalized roots of $(W,S)$, denoted by $\widehat{\Phi}$, is the intersection of $\bigcup_{\beta\in\Phi}\mathbb{R}\beta$ with $P$. For any $\beta=\sum_{i=1}^{n}k_{i}\alpha_{i}\in\Phi^{+}$, there is a unique normalized root $\widehat{\beta}=\frac{\beta}{\sum_{i=1}^{n}k_{i}}\in\widehat{\Phi}$ associated with it. The barycentric coordinates of $\beta$ are given by
$(\frac{k_{1}}{\sum_{i=1}^{n}k_{i}},\frac{k_{2}}{\sum_{i=1}^{n}k_{i}},\cdots,\frac{k_{n}}{\sum_{i=1}^{n}k_{i}}).$
@@ 2.6. Lexicographic reflection order
Let $V=\mathbb{R}\Phi$ and let the root basis $\Pi=\{\alpha_{1},\alpha_{2},\dots,\alpha_{n}\}$. Choose an ordered basis $\{v_{1},v_{2},\cdots,v_{n}\}$ of $V$. If the root basis is linearly independent, a frequent choice of this basis is $v_{i}=\alpha_{i},1\leq i\leq n$ , but we do not assume this. Each root can be written uniquely as $\sum_{i=1}^{n}k_{i}v_{i}.$ This induces a lexicographic ordering on $\mathbb{R}^{n}$:
@@ 2.8. Limit root and isotropic cone
The normalized isotropic cone is the intersection of the standard affine hyperplane with the isotropic cone. An accumulation point of $\widehat{\Phi}$ (the set of projective roots) is called a limit root. It has been proved in [9] that all limit roots lie on the normalized isotropic cone. The concept of limit roots was recently introduced as a tool for better visualizing the positive roots. This has proved instrumental in investigating the distribution of roots in infinite Coxeter systems of low rank, providing deeper insight into concepts such as the imaginary cone, biclosed sets, the dominance order, and reflection subgroups. For this notion and its application, see [3], [5], [7], [9], [8], [10] and [12].
@@
Figure 1. The normalized roots and the normalized isotropic cone of a rank 3 universal Coxeter system.
@@ 3. Dense Subset
Throughout this section, let $(W,S)$ be a rank 3 universal Coxeter system, i.e. $S=\{s_{1},s_{2},s_{3}\}$ and
@@
Let the root system $\Phi$ be constructed via the standard reflection representation as described in Subsection 2.2. Impose a total order on $\Pi$ by requiring $\alpha_{s_{1}}\prec\alpha_{s_{2}}\prec\alpha_{s_{3}}$. Denote the associated lexicographic reflection order by $\prec_{\mathrm{reflex}}$. This section is devoted to the analysis of the order type of this particular reflection order. For convenience, we write $\alpha_{i}$ for $\alpha_{s_{i}}$.
@@
is a positive root. Therefore, $c_{1}+2c_{2}+2c_{3}\geq 0$, and consequently, $c_{2}+c_{3}\geq\frac{c_{1}}{2}$. If $c_{1}>0$, then
@@
Assume that $c\in(0,\frac{2}{3})$, and that the set
@@
is infinite. Furthermore, this set consists precisely of the positive roots of a maximal infinite dihedral reflection subgroup.
@@
contains at least two elements. There exists a unique maximal dihedral reflection subgroup $W^{\prime}$ whose set of positive roots $\Phi_{W^{\prime}}^{+}$ contains these two roots. Furthermore, $\Phi_{W^{\prime}}^{+}$ consists precisely of those positive roots whose first barycentric coordinate is $c$, and the normalized roots of them lie on a line segment parallel to the side connecting $\widehat{\alpha_{2}}$ and $\widehat{\alpha_{3}}$ of the equilateral triangle in Figure 1 (as depicted by the green line segment there). Since $c<\frac{2}{3}$, this line segment intersects with the normalized isotropic cone (the circle inscribed in the triangle). By Proposition 1.5(ii) in [9] this implies that $W^{\prime}$ is an infinite dihedral reflection subgroup, and therefore $\Phi_{W^{\prime}}^{+}$ is also infinite. Hence, there are an infinite number of positive roots with $\frac{c_{1}}{c_{1}+c_{2}+c_{3}}=c$.
If otherwise $c_{2}=c_{3}=k$, we have $B(c_{1}\alpha_{1}+k\alpha_{2}+k\alpha_{3},c_{1}\alpha_{1}+k\alpha_{2}+k\alpha_{3})=1.$ Therefore, $c_{1}^{2}-4c_{1}k=c_{1}(c_{1}-4k)=1$. Since $c_{1}$ and $c_{1}-4k$ are both integers, $c_{1}=1,k=0$, and thus $c_{1}\not\in(0,\frac{2}{3})$. Therefore, the assertion of the Lemma follows. ∎
@@
The set $\{\frac{c_{1}}{c_{1}+c_{2}+c_{3}}\mid\,c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\}\cap(0,\frac{2}{3})$ is infinite and is dense under the natural order.
@@
The set of positive roots of the parabolic subgroup generated by $s_{1}$ and $s_{3}$ contain infinitely many roots whose first barycentric coordinate is between 0 and $\frac{2}{3}$. Therefore this set is infinite. Let $x,y\in\left(0,\tfrac{2}{3}\right)$ with $x\neq y$ be such that there exist positive roots $\alpha^{\prime}$ and $\beta^{\prime}$ whose first barycentric coordinates are $x$ and $y$, respectively. By Lemma 3.2, there exists a maximal infinite dihedral reflection subgroup $W^{\prime}$ (resp. $W^{\prime\prime}$) whose set of positive roots contains $\alpha$ (resp. $\beta$). Furthermore, any positive root of $W^{\prime}$ (resp. $W^{\prime\prime}$) has its first barycentric coordinate equal to $x$ (resp. $y$). Since the limit roots of $W^{\prime}$ and $W^{\prime\prime}$ lie on the (normalized) isotropic cone, two roots $\alpha^{\prime}\in\Phi_{W^{\prime}}^{+},\beta^{\prime}\in\Phi_{W^{\prime\prime}}^{+}$ can be chosen such that $\widehat{\alpha^{\prime}}$ and $\widehat{\beta^{\prime}}$ are close enough to the normalized isotropic cone for the line segment connecting them to cut the normalized isotropic cone. See Figure 2 for an illustration of one possible situation. Then, by Proposition 1.5(ii) in [9], the maximal dihedral reflection subgroup containing the reflections in $\alpha^{\prime}$ and $\beta^{\prime}$ is infinite, and there exists a positive root $\gamma$ in the set of positive roots of this maximal dihedral reflection subgroup such that $\gamma$ is a positive linear combination of $\alpha^{\prime}$ and $\beta^{\prime}$. Therefore, he first barycentric coordinate of $\gamma$ is necessarily between $x$ and $y$. ∎
@@
It is natural to ask whether there is a natural or intrinsic characterization of the set
$\left\{\frac{c_{1}}{c_{1}+c_{2}+c_{3}}\,\middle|\,c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\right\}\cap\left(0,\tfrac{2}{3}\right)?$
@@
If $\frac{c_{1}}{c_{1}+c_{2}+c_{3}}=0,$ then $\alpha:=c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}$ is a root contained in the root system of the (infinite dihedral) parabolic subgroup generated by $s_{2},s_{3}$. These roots form an initial interval of $\prec_{\mathrm{reflex}}.$ The corresponding normalized roots lie on the side connecting $\widehat{\alpha_{2}}$ and $\widehat{\alpha_{3}}$ in the equilateral triangle in Figure 1. Write $\Phi^{+}_{1}=\{c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\mid\,c_{1}=0\}.$
If $\frac{c_{1}}{c_{1}+c_{2}+c_{3}}=\frac{2}{3}$, then $\alpha:=c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}$ is a root contained in the root system of the infinite maximal dihedral reflection subgroup generated by $s_{2\alpha_{1}+\alpha_{2}},s_{2\alpha_{1}+\alpha_{3}}$. The corresponding normalized roots lie on the blue line segment in Figure 1. Write $\Phi^{+}_{3}=\{c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\mid\,\frac{c_{1}}{c_{1}+c_{2}+c_{3}}=\frac{2}{3}\}$ and write $\Phi^{+}_{2}=\{c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\mid\,\frac{c_{1}}{c_{1}+c_{2}+c_{3}}\in(0,\frac{2}{3})\}.$
@@
Note that by the previous analysis, the restriction of $\prec_{\mathrm{reflex}}$ to $\Phi^{+}_{1}$ (resp. $\Phi^{+}_{3}$) has the order type $\omega+\omega^{*}$.
Let $c^{\prime}\in\{\frac{c_{1}}{c_{1}+c_{2}+c_{3}}\mid\,c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\}\cap(0,\frac{2}{3})$. By Lemma 3.2, the roots in the set $C^{\prime}:=\{c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\mid\frac{c_{1}}{c_{1}+c_{2}+c_{3}}=c^{\prime}\}$ are the positive roots of an infinite dihedral subgroup. Therefore, by the definition of the lexicographic reflection order and the fact that $\omega+\omega^{*}$ is the only possible order type of a reflection order on a dihedral group, we conclude that the roots in $C^{\prime}$ form an interval $I_{c^{\prime}}$ in $\prec_{\mathrm{reflex}}$ of order type $\omega+\omega^{*}$. By the definition of lexicographic reflection order, one has $I_{c^{\prime}}\prec I_{c^{\prime\prime}}$ if and only if $c^{\prime}<c^{\prime\prime}$. On the other hand, $\Phi^{+}_{2}$ is the disjoint union of $I_{c^{\prime}},c^{\prime}\in\{\frac{c_{1}}{c_{1}+c_{2}+c_{3}}\mid\,c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\}\cap(0,\frac{2}{3})$, which is dense and is infinite by Lemma 3.3. Therefore, by Theorem 2.8 of [13], we can conclude that the natural total order on
$\left\{\frac{c_{1}}{c_{1}+c_{2}+c_{3}}\,\middle|\,c_{1}\alpha_{1}+c_{2}\alpha_{2}+c_{3}\alpha_{3}\in\Phi^{+}\right\}\cap\left(0,\tfrac{2}{3}\right)$
has the order type $\eta$, the order type of the rational numbers. Consequently, the restriction of $\prec_{\mathrm{reflex}}$ to $\Phi_{2}^{+}$ has order type $(\omega+\omega^{*})\eta$.
@@
We record the following two Lemmas for later use.
@@
is not empty. By symmetry, Lemma 3.2 guarantees that this set
@@
is precisely the set of the positive roots of a maximal infinite dihedral reflection subgroup $U^{\prime}$. The normalized roots of them lie on a line segment parallel to the side connecting $\widehat{\alpha_{1}}$ and $\widehat{\alpha_{3}}$ of the equilateral triangle in Figure 1 (as depicted by the purple line segment there). The first barycentric coordinates of the positive roots of $\Phi_{U^{\prime}}^{+}$ vary between 0 and $\frac{2}{3}$. If $\alpha,\beta\in\Phi_{U^{\prime}}^{+}$ are two roots such that the first barycentric coordinate of $\alpha$ is smaller than that of $\beta$, then $\alpha\prec_{\mathrm{reflex}}\beta.$ Then, by Lemma 3.3, the interval $(\alpha,\beta)$ is infinite, since there are infinitely many roots whose first barycentric coordinates lie strictly between those of $\alpha$ and $\beta$. ∎
Remark 3.7.
Let $\alpha$ and $\beta$ be two distinct positive roots. It is not hard to see that, by suitably choosing a basis of $\mathbb{R}\Phi_{W}$, one can always find a lexicographic reflection order $\prec$ such that the interval $\alpha,\beta]$ coincides with the restriction of $\prec$ to the maximal reflection subgroup containing $\alpha$ and $\beta$. In particular, $\alpha,\beta]$ is scattered.
Lemma 3.8.
Let $\prec$ be any reflection order on $(W,S)$, and let $k$ be a positive integer. There exists a subset $E\subset\Phi^{+}$ such that the restriction of $\prec$ to $E$ has order type $(\omega+\omega^{*})[k]$.
Proof.
One can always find $k+1$ normalized roots $\widehat{\alpha}_{1},\widehat{\alpha}_{2},\dots,\widehat{\alpha}_{k+1}$ sufficiently close to the isotropic cone such that the line segment joining any two of them intersects the normalized isotropic cone. Without loss of generality, assume that
$\widehat{\alpha}_{1}\prec\widehat{\alpha}_{2}\prec\cdots\prec\widehat{\alpha}_{k+1}.$
For each $i$, write $\alpha_{i},\alpha_{i+1}]$ for the interval between $\alpha_{i}$ and $\alpha_{i+1}$ with respect to $\prec$.
Let $U_{i}$ denote the intersection of $\alpha_{i},\alpha_{i+1}]$ with the positive system $\Phi^{+}_{W_{\{s_{\alpha_{i}},s_{\alpha_{i+1}}\}}}$ of the dihedral reflection subgroup $W_{\{s_{\alpha_{i}},s_{\alpha_{i+1}}\}}$, and set
$E=\bigcup_{i=1}^{k}U_{i}.$
Since the line segment joining $\alpha_{i}$ and $\alpha_{i+1}$ intersects the normalized isotropic cone, the restriction of $\prec$ to $U_{i}$ has order type $\omega+\omega^{*}$. Consequently, the restriction of $\prec$ to $E$ has the desired order type. ∎